impedance of (2 + j4)n per phase, feeds two balanced 2.7 A three-phase line, which has an three-phase loads that are connectecan I have the correct solutions, answers are given
Answer #1

-24 7 349 160 tivo) B per phase equivalent ckt Cek (by converting a convertinga Anconnected lead into its equivalent star conze = 22-36 Z-10.305 = 29-34 -Zstzq = 245$ +22-24 = 24 presies In = Slime : Vfhca) Zgt ze 1201 24 I 5 * pour le real power delViz-lead = 19364 219.095 v c) cureat imper phase in dar conected load (5) I Vph (lead) ZA (volela 111.8 2-10.305 14.8410.305_load 3 ( Vectoad * IXC8 (19:595-56-565) ahsashed = .3x193164 x 2.581 x COS(-86-87) 1199.48 W Othere 3 (Vil= load XI, X sin 6